AnimeDude743:

x + y = 7 x + y = 6 What is the solution of the system shown? (6, 12) ( , 14) (10 , 5)

astroo:

66

Florisalreadytaken:

\[\begin{cases}x + y = 7\\ x + y = 6 \end{cases}\] from the first equation we can write that: \( x + y = 7 \Rightarrow x=7-y \) lets plug that information in the second equation: \[ x + y = 6 \] \[ 7-y + y = 6 \] \[ 7=6 \] \[ 0=6-7 \] \[ 0=-1 \] this has got no solution...

Shayla:

@florisalreadytaken wrote:
\[\begin{cases}x + y = 7\\ x + y = 6 \end{cases}\] from the first equation we can write that: \( x + y = 7 \Rightarrow x=7-y \) lets plug that information in the second equation: \[ x + y = 6 \] \[ 7-y + y = 6 \] \[ 7=6 \] \[ 0=6-7 \] \[ 0=-1 \] this has got no solution...
Florrr I need your help with something, I'd message you and ask but ig you've got messaging turned off 🥺